Radius 13 inch, height 34 inch find its volume
Answer by FrankM(1040)

If this is a cylinder, the volume is base times height, or
pi*r^2*h = 3.1416*169*34=18051.6336 cubic inches.
If this was a cone, please divide by 3. A cone's volume is 1/3 the volume of its enclosing cylinder.